OpenStudy (anonymous):

simplify 9 + 4b + c - 9b - 1

OpenStudy (anonymous):

\[\LARGE ax+bx=(a+b)x\] use this..

OpenStudy (anonymous):

\[\LARGE 4b-9b=?b\]

OpenStudy (anonymous):

can you do it...?

OpenStudy (anonymous):

I think so. But after I do this formula, how does "c" and 9 fit in? does 9 or c go first, and I dont understand whether or not I shoul put "+" or "-"

OpenStudy (anonymous):

\[\LARGE 9+4b+c-9b-1=(4-9)b+c+(9-1)\] can you do it now ?? c and 9 doesn't fit , leave them separated ...

OpenStudy (anonymous):

Oh so -5b goes in between them?

OpenStudy (anonymous):

good Question. :)

OpenStudy (anonymous):

.... what's your final answer? (Even if it's wrong it doesn't matter, I want to see what you've done so far ;) )

OpenStudy (anonymous):

9 + -5b -C

OpenStudy (anonymous):

hmmm ... (9-1)=? -5b it's ok, well done. just take off plus before minus , and c should be positive... tell me what you get?

OpenStudy (anonymous):

wait...where does the 1 come from? O_O (and I meant " - c" forgot the space)

OpenStudy (anonymous):

craftshark98 simplify 9 + 4b + c - 9b - 1 this is what you wrote... I SEE a -1 there .. do you see it ?

OpenStudy (anonymous):

this is the final answer ... \[\huge -5b+c+8\]

OpenStudy (anonymous):

oh nevermind duh theres a one at the end x] and thanks! I just had to think of it like a list of numbers instead of an equation...